# Low-Altitude ISAC with Rotatable Active and Passive Arrays

### v1

@@ III Proposed Algorithm
In this section, we convert (P1) to a penalized form and solve it by resorting to AO technique across three blocks: BS transmit beamforming $\boldsymbol{W}$, RIS passive beamforming $\boldsymbol{\theta}$, and array rotations $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$. The three subproblems are handled by a quadratic transform (QT) and a majorization-minimization (MM) step for $\boldsymbol{W}$, Riemannian optimization for $\boldsymbol{\theta}$, and the projection gradient-based method for $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, respectively.
III-A Penalty Method and AO Framework
The beam pattern similarity constraints (22b), with a high-order non-convex MSE expression, limit the feasible region and hinder the solution of (P1). To handle this difficulty, we incorporate (22b) into the objective function as a hinge penalty term, obtaining a more tractable reformulated problem
$\text{(P1.1)}:\underset{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}}{\max}$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}\left(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}},\boldsymbol{\theta},\boldsymbol{W}\right)\right)}$
$\!\!\!\!-\rho\max\!\left\{\text{MSE}\left(\boldsymbol{r}^{\text{B}}\!,\boldsymbol{r}^{\text{R}}\!,\boldsymbol{\theta}\!,\boldsymbol{W}\right)\!-\!\eta_{\text{th}},0\right\}\!\!\!$ (23)
s.t. $\left(22\text{c}\right),\left(22\text{d}\right),\left(22\text{e}\right),\left(22\text{f}\right)$
where the adjustable penalty parameter $\rho\geq 0$ penalizes violations of (22b) with the hinge penalty function $\max\!\left\{\text{MSE}\!-\!\eta_{\text{th}},0\right\}$. When $\rho\rightarrow\infty$, any solution $\{(\boldsymbol{r}^{\text{B}})^{*},(\boldsymbol{r}^{\text{R}})^{*},\boldsymbol{\theta}^{*},\boldsymbol{W}^{*}\}$ of (P1.1) should meet the condition $\text{MSE}\leq\eta_{\text{th}}$, establishing the equivalence between (P1) and (P1.1). Since setting $\rho$ large at the beginning renders this approach ineffective, we initialize $\rho$ by a small value to find a good starting point, solve problem (P1.1) iteratively with gradually increased $\rho$, until $\text{MSE}\leq\eta_{\text{th}}$ holds and the objective function reaches the convergence criteria. $\rho$ increases only when $\mathrm{MSE}>\eta_{\text{th}}$ persists; once the constraint is satisfied, $\rho$ is kept fixed to avoid unnecessary rate loss.
Even with fixed $\rho$, due to the non-concave objective function and the non-convex constraints with closely coupled variables, (P1.1) remains highly non-convex and challenging to solve directly. Nevertheless, observing that the problem becomes more tractable when optimizing one variable block at a time, we resort to the AO method: at each outer iteration, we cyclically optimize the transmit beamforming $\boldsymbol{W}$, the RIS phase shifts $\boldsymbol{\theta}$, and the array rotations $\{\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}}\}$, while keeping the others fixed, as detailed in the following subsections.
@@ III-B QT- and MM-based Transmit Beamforming
With fixed $(\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, the subproblem optimizing $\boldsymbol{W}$ reads
$\text{(P2)}:\underset{\boldsymbol{W}}{\max}\,\,$ $\sum_{k=1}^{K}{\log_{2}\left(1+\text{SINR}_{k}(\boldsymbol{W})\right)}$
$-\rho\max\left\{\text{MSE}(\boldsymbol{W})-\eta_{\text{th}},0\right\}$ (24)
s.t. $\left(22\text{c}\right),$
where the SINR and MSE expressions reduce to
$\text{SINR}_{k}\left(\boldsymbol{W}\right)=\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1,i\neq k}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}},$ (25a)
$\text{MSE}\left(\boldsymbol{W}\right)=\frac{1}{A}\sum_{a=1}^{A}{\left|\|\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\|_{2}^{2}-\mathcal{P}_{\text{d},a}\right|^{2}}.$ (25b)
The power constraint (22c) is convex, and the non-concavity of (24) originates from the fractional SINR and fourth-order MSE expressions, which are sequentially tackled as follows.
@@ III-B1 Sensing MSE term with MM-based second-order upper bound
$g_{a}\left(\boldsymbol{W}\right)$ $\triangleq\big(\lVert\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\rVert_{2}^{2}-\mathcal{P}_{\text{d},a}\big)^{2},$ (26a)
$=\big(\text{Tr}(\boldsymbol{W}^{H}\boldsymbol{S}_{a}\boldsymbol{W})-\mathcal{P}_{\text{d},a}\big)^{2},$ (26b)
@@
$=4\big(\text{Tr}(\boldsymbol{W}^{H}\boldsymbol{S}_{a}\boldsymbol{W})-\mathcal{P}_{\text{d},a}\big)\boldsymbol{S}_{a}\boldsymbol{W}.$ (28b)
@@
where $L_{a}^{\text{lip}}$ is any Lipschitz constant on the Frobenius ball $\{\boldsymbol{W}:\|\boldsymbol{W}\|_{F}^{2}\leq P_{\mathrm{B}}\}$ for $\nabla g_{a}(\boldsymbol{W})$. Using matrix-norm inequalities, one obtains a choice for a global Lipschitz constant
$L_{a}^{\text{lip}}=12P_{\text{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}^{2}\,+4\mathcal{P}_{\text{d},a}\lVert\boldsymbol{S}_{a}\rVert_{2},$ (30)
with $\|\boldsymbol{S}_{a}\|_{2}=\|\boldsymbol{f}_{\mathrm{S},a}\|_{2}^{2}$.
Proof: For any two points $\boldsymbol{W}_{1}$ and $\boldsymbol{W}_{2}$ on $g_{a}(\boldsymbol{W})$, we have
@@
$=4\big\|(\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathcal{P}_{\text{d},a})\boldsymbol{S}_{a}(\boldsymbol{W}_{1}-\boldsymbol{W}_{2})$
$\quad\quad+(\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2}))\boldsymbol{S}_{a}\boldsymbol{W}_{2}\big\|$ (31a)
$\leq 4\lVert\boldsymbol{S}_{a}\rVert_{2}\big(|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathcal{P}_{\text{d},a}|\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}$
$\quad\quad+|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})\!-\!\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2})|\lVert\boldsymbol{W}_{2}\rVert_{F}\big).\!\!\!$ (31b)
Then, according to the following inequalities for terms in (31b)
$|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1}t)-\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2})|$
$=|\text{Tr}((\boldsymbol{W}_{1}-\boldsymbol{W}_{2})^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\text{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}(\boldsymbol{W}_{1}-\boldsymbol{W}_{2}))|$
@@
$|\text{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathcal{P}_{\text{d},a}|\leq\lVert\boldsymbol{S}_{a}\rVert_{2}\lVert\boldsymbol{W}_{1}\rVert_{F}^{2}+\mathcal{P}_{\text{d},a},$ (33)
$\lVert\boldsymbol{W}_{1}\rVert_{F},\lVert\boldsymbol{W}_{2}\rVert_{F}\leq\sqrt{P_{B}},$ (34)
we substitute (32)-(34) into (31) and finally obtain
@@
$\leq 4\lVert\boldsymbol{S}_{a}\rVert_{2}\left(P_{B}\lVert\boldsymbol{S}_{a}\rVert_{2}+\mathcal{P}_{\text{d},a}\right)\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}$
$\quad+4\lVert\boldsymbol{S}_{a}\rVert_{2}(2\sqrt{P_{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F})\sqrt{P_{B}}$
$=\left(12P_{B}\lVert\boldsymbol{S}_{a}\rVert_{2}^{2}+4\lVert\boldsymbol{S}_{a}\rVert_{2}\mathcal{P}_{\text{d},a}\right)\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F},$ (35)
which is exactly the definition of the Lipschitz constant.    $\blacksquare$
Consequently, replacing each $g_{a}(\boldsymbol{W})$ by its upper bound yields a concave penalization term $\rho\max\{\mathcal{U}_{a}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)})-\!\eta_{\text{th}},0\}$. At each iteration, (29) provides a tight lower bound on $\boldsymbol{W}^{(t)}$ to maximize the objective function (24), guaranteeing non-decreasing penalized objective values in updates to $\boldsymbol{W}$ with a standard MM step.
III-B2 Communication SINR term with Lagrangian dual transform (LDT) and QT for fractions
@@
Next, introducing complex-valued variables $\{y_{k},\forall k\}$, we apply QT to transform the fractional SINR term into
@@
$=\underset{y_{k}\in\mathbb{C}}{\max}\,\,2\text{Re}\{y_{k}^{*}\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\}\!-\!\left|y_{k}\right|^{2}\Big(\!\sum_{i=1}^{K+M}\!{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}\!+\!\sigma_{k}^{2}\Big).\!\!$ (37)
@@
$y_{k}^{*}=\frac{\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}}{\sum_{i=1}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}}.$ (39)
@@
$\mathcal{Y}\left(\boldsymbol{W}\mid\mu_{k},y_{k}\right)$
$\!=\!\frac{1\!\!+\!\mu_{k}}{\ln 2}\Big(2\text{Re}\{y_{k}^{*}\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\}\!-\!\left|y_{k}\right|^{2}\!\Big(\!\sum_{i=1}^{K+M}{\!\!}|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}\!\!+\!\sigma_{k}^{2}\Big)\!\Big).\!\!\!$ (40)

### v2

@@ III Proposed Algorithm
In this section, we develop an efficient algorithm to tackle the non-concave objective and the non-convex constraints in (P1), where the optimization variables are tightly coupled through both the multiuser SINRs and the sensing NMSE.
III-A AO framework
A key observation is that (P1) becomes considerably more tractable when optimizing one variable block at a time. Hence, we resort to an AO framework that alternately updates three blocks in an iterative manner: the BS transmit beamforming matrix $\boldsymbol{W}$, the RIS passive beamforming vector $\boldsymbol{\theta}$, and the array rotations $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$.
Importantly, the sensing metric is $\text{NMSE}(\iota)$ in (22), where the scaling parameter $\iota$ is introduced to decouple global amplitude mismatch from angular shape mismatch. We update $\iota$ once per AO outer iteration using the closed-form minimizer $\iota^{\star}$ in (23), and keep it fixed when solving the three inner subproblems for $\boldsymbol{W}$, $\boldsymbol{\theta}$, and $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$. This outer update is computationally negligible and does not introduce additional constraints. Moreover, since $\iota^{\star}$ minimizes $\text{NMSE}(\iota)$ for fixed $(\boldsymbol{W},\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$, and the sum-rate term is independent of $\iota$, updating $\iota$ with $\iota^{\star}$ makes the objective value in (24a) non-decreasing. Therefore, the $\iota$-update does not undermine the monotonic improvement of the overall AO procedure; it simply refines the sensing metric by removing the best-fit scaling mismatch at the current iterate.
Then, in each iteration, the resulting three subproblems are handled by: i) a QT combined with a MM step for updating $\boldsymbol{W}$, ii) a Riemannian optimization method over the complex circle manifold for updating $\boldsymbol{\theta}$, and iii) a projected gradient-based method for updating $(\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ under box constraints, respectively. Due to the intrinsic non-convexity of (P1), global optimality is generally intractable. The proposed AO-based algorithm aims at computing a locally optimal or suboptimal solution that achieves a favorable communication-sensing trade-off. In practice, such a suboptimal solution is meaningful because it is obtained with low complexity, respects all hardware constraints such as unit-modulus phases and mechanical rotation limits, and can be efficiently warm-started and re-optimized as the low-altitude geometry evolves.
@@ III-B QT- and MM-based Transmit Beamforming
With $(\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$ and $\iota$ fixed, the subproblem for optimizing $\boldsymbol{W}$ becomes
$\text{(P2)}:\underset{\boldsymbol{W}}{\max}\,$ $\sum_{k=1}^{K}\log_{2}\left(1+\mathrm{SINR}_{k}(\boldsymbol{W})\right)-\rho\text{NMSE}(\boldsymbol{W})$ (25)
[…]
and the SINR and NMSE expressions reduce to
$\text{SINR}_{k}(\boldsymbol{W})$ $=\frac{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}|^{2}}{\sum_{i=1,i\neq k}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}},$ (26a)
$\text{NMSE}(\boldsymbol{W})$ $=\frac{\sum_{a=1}^{A}\big|\|\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\|_{2}^{2}-\iota\mathcal{P}_{d,a}\big|^{2}}{\sum_{a=1}^{A}\left|\iota\mathcal{P}_{d,a}\right|^{2}}$ (26b)
$=\frac{1}{D}\sum_{a=1}^{A}\big|\|\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\|_{2}^{2}-\bar{\mathcal{P}}_{d,a}\big|^{2},$ (26c)
where we denote $D\triangleq\sum_{a=1}^{A}\left|\iota\mathcal{P}_{d,a}\right|^{2}$ and $\bar{\mathcal{P}}_{d,a}\triangleq\iota\mathcal{P}_{d,a}$ determined by the fixed $\iota$ for brevity. Here, $\boldsymbol{f}_{k}$ denotes the effective communication channel of user $k$, and $\boldsymbol{f}_{\text{S},a}$ denotes the effective sensing steering vector at grid point $a$, both determined by the fixed $(\boldsymbol{\theta},\boldsymbol{r}^{\text{B}},\boldsymbol{r}^{\text{R}})$. As the power constraint (24b) is convex, the non-concavity of (25) mainly stems from the fractional SINR terms in (26a) and the fourth-order dependence of the sensing loss $\|\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\|_{2}^{2}$ in (26c). In the sequel, we first apply the QT to reformulate the sum-rate into a tractable form, and then employ an MM step to construct a tight surrogate for the sensing-related quartic term, leading to a closed-form update of $\boldsymbol{W}$ with bisection search.
@@ III-B1 Sensing MSE term with MM-based second-order upper bound
$g_{a}(\boldsymbol{W})$ $\triangleq\big(\lVert\boldsymbol{W}^{H}\boldsymbol{f}_{\text{S},a}\rVert_{2}^{2}-\bar{\mathcal{P}}_{\text{d},a}\big)^{2},$ (27a)
$=\big(\text{Tr}(\boldsymbol{W}^{H}\boldsymbol{S}_{a}\boldsymbol{W})-\bar{\mathcal{P}}_{\text{d},a}\big)^{2},$ (27b)
@@
$=4\big(\mathrm{Tr}(\boldsymbol{W}^{H}\boldsymbol{S}_{a}\boldsymbol{W})-\bar{\mathcal{P}}_{\text{d},a}\big)\boldsymbol{S}_{a}\boldsymbol{W}.$ (29b)
@@
where $L_{a}^{\text{lip}}$ is any Lipschitz constant on the Frobenius ball $\{\boldsymbol{W}:\|\boldsymbol{W}\|_{F}^{2}\leq P_{\text{B}}\}$ for $\nabla g_{a}(\boldsymbol{W})$. Using matrix-norm inequalities, one obtains a choice for a global Lipschitz constant
$L_{a}^{\text{lip}}=12P_{\text{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}^{2}+4\bar{\mathcal{P}}_{\text{d},a}\lVert\boldsymbol{S}_{a}\rVert_{2},$ (31)
with $\|\boldsymbol{S}_{a}\|_{2}=\|\boldsymbol{f}_{\text{S},a}\|_{2}^{2}$.
Proof:
For any two points $\boldsymbol{W}_{1}$ and $\boldsymbol{W}_{2}$ on the function $g_{a}(\boldsymbol{W})$, according to matrix norm inequalities, we have
@@
$=4\big\|(\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\bar{\mathcal{P}}_{\text{d},a})\boldsymbol{S}_{a}(\boldsymbol{W}_{1}-\boldsymbol{W}_{2})$
$\quad+(\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathrm{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2}))\boldsymbol{S}_{a}\boldsymbol{W}_{2}\big\|$
$\leq 4\lVert\boldsymbol{S}_{a}\rVert_{2}\big(|\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\bar{\mathcal{P}}_{\text{d},a}|\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}$
$\quad+|\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})\!-\!\mathrm{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2})|\lVert\boldsymbol{W}_{2}\rVert_{F}\big).\!$ (32)
Then, we construct the following inequalities for terms related to traces in (III-B1)
$|\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathrm{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{2})|$
$=|\mathrm{Tr}((\boldsymbol{W}_{1}-\boldsymbol{W}_{2})^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\mathrm{Tr}(\boldsymbol{W}_{2}^{H}\boldsymbol{S}_{a}(\boldsymbol{W}_{1}-\boldsymbol{W}_{2}))|$
@@
$|\mathrm{Tr}(\boldsymbol{W}_{1}^{H}\boldsymbol{S}_{a}\boldsymbol{W}_{1})-\bar{\mathcal{P}}_{\text{d},a}|\leq\lVert\boldsymbol{S}_{a}\rVert_{2}\lVert\boldsymbol{W}_{1}\rVert_{F}^{2}+\bar{\mathcal{P}}_{\text{d},a},$ (34)
where the power limitation holds
$\lVert\boldsymbol{W}_{1}\rVert_{F}\leq\sqrt{P_{\text{B}}},\quad\lVert\boldsymbol{W}_{2}\rVert_{F}\leq\sqrt{P_{\text{B}}}.$ (35)
Substituting (33), (34), and (35) into (III-B1), we obtain
@@
$\leq 4\lVert\boldsymbol{S}_{a}\rVert_{2}\left(P_{\text{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}+\bar{\mathcal{P}}_{\text{d},a}\right)\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F}$
$\quad+4\lVert\boldsymbol{S}_{a}\rVert_{2}(2\sqrt{P_{\text{B}}}\lVert\boldsymbol{S}_{a}\rVert_{2}\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F})\sqrt{P_{\text{B}}}$
$=\left(12P_{\text{B}}\lVert\boldsymbol{S}_{a}\rVert_{2}^{2}+4\lVert\boldsymbol{S}_{a}\rVert_{2}\bar{\mathcal{P}}_{\text{d},a}\right)\lVert\boldsymbol{W}_{1}-\boldsymbol{W}_{2}\rVert_{F},$ (36)
which is exactly the definition of the Lipschitz constant. ∎
Consequently, at each iteration, replacing each $g_{a}(\boldsymbol{W})$ by its concave upper bound $\mathcal{U}_{a}(\boldsymbol{W}\mid\boldsymbol{W}^{(t)})$ on $\boldsymbol{W}^{(t)}$ in (30) provides a tight lower bound to maximize the objective function $\rho\text{NMSE}(\boldsymbol{W})$ of (25), guaranteeing non-decreasing objective values in updates to $\boldsymbol{W}$ with a standard $M$ step.
III-B2 Communication SINR term with LDT and QT
@@
Next, introducing complex-valued variables $\{\eta_{k},\forall k\}$, we apply QT to transform the fractional SINR term into
@@
$=\underset{\eta_{k}\in\mathbb{C}}{\max}\,\,2\text{Re}\{\eta_{k}^{*}\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\}\!-\!\left|\eta_{k}\right|^{2}\Big(\!\sum_{i=1}^{K+M}\!{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}\!+\!\sigma_{k}^{2}\Big).\!\!$ (38)
@@
$\eta_{k}^{*}=\frac{\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}}{\sum_{i=1}^{K+M}{|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}}+\sigma_{k}^{2}}.$ (40)
@@
$\mathcal{Y}\left(\boldsymbol{W}\mid\mu_{k},\eta_{k}\right)$
$\!=\!\frac{1\!\!+\!\mu_{k}}{\ln 2}\Big(2\text{Re}\{\eta_{k}^{*}\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{k}\}\!-\!\left|\eta_{k}\right|^{2}\!\Big(\!\sum_{i=1}^{K+M}{\!\!}|\boldsymbol{f}_{k}^{H}\boldsymbol{w}_{i}|^{2}\!\!+\!\sigma_{k}^{2}\Big)\!\Big).\!\!\!$ (41)